Find the corresponding area between 𝑧 = 0 and each of the following.

1. z = 0.85

2. z = 1.27

3. z = 2.86

4. z = βˆ’1.05

5. z = βˆ’2.96


1
Expert's answer
2021-03-31T13:45:04-0400

s = |f(z) - 0.5|

f(z) is taken from a standart normal distribution table:



a) 0.80234 - 0.5 = 0.30234

b) 0.89796Β - 0.5 = 0.39796

c) 0.99788 - 0.5 = 0.49788

d) |0.14686 - 0.5| = 0.35314

e) |0.00154 - 0.5| = 0.49846